# A Literature Review of Literature Reviews in Pattern Analysis and Machine Intelligence

### v5

@@ 5.1 Appraising Reviews across Various Topics
A fundamental need may lie in the assessment of reviews across various fields, as most reviews are primarily aimed at readers who are relatively new to a field and may lack an in-depth understanding of the field. For instance, a researcher working in a popular field like object detection may start exploring a less familiar subfield such as optical character recognition (a subfield of computer vision). When encountering a review with only 100 citations, the researcher might dismiss its importance, unaware that in a niche area, 100 citations can actually indicate significant influence. This could lead to an oversight of critical insights in the subfield, simply due to differing citation norms between areas. The proposed $TNCSI$ offers an effective solution by automatically identifying the field of a given article and dynamically estimating its normalized impact based on the field and its current citation count. Illustrative examples are presented in Tab. IV.
TABLE IV: Example Use Cases of $TNCSI$: A similar number of citations (Cites) in various topics may lead to notably different $TNCSI$ values.
Title Cites Topic $TNCSI$
@@
5.2 Tracking Citation Momentum with IEI
Although $TNCSI$ enables cross-domain comparisons, it fails to track changes in citation trends. In such cases, the $IEI$ can offer additional insight. A higher $IEI$ suggests that the paper’s rate of gaining new citations has been increasing over time, reflecting a growing level of recognition within the academic community. Conversely, an $IEI$ below zero implies that the paper’s citation rate has been declining recently. Table V provides examples of $IEI$ applications. In some cases, selecting a paper with a higher $IEI$ may be preferable to choosing one with a higher $TNCSI$, as it reflects a more recent consensus among researchers.
TABLE V: Example Use Cases of $IEI$: A positive $IEI$ indicates that the paper is gaining increasing attention.
Title Cites $TNCSI$ $IEI$
@@ 5.3 The Role of Reference Quality in Selecting Reviews
In Sec. 4.3, we briefly discuss the importance of references, noting that the quality of references shows a moderately strong and significant positive correlation with the scholarly impact of a review. Thus, $RQM$ functions as an intuitive metric, allowing researchers to efficiently gauge the average quality of references within a review. Based on the statistical patterns shown in Fig. 6, reviews with an $RQM$ greater than 0.8, an $ARQ$ within the range of 0.7 to 0.9, and an $S_{MP}$ not exceeding 10 are more likely to achieve greater academic impact. Examples of $RQM$ can be found in Tab. VI.
TABLE VI: Example Use Cases of $RQM$: A higher $RQM$ reflects higher scholarly impact ($ARQ$) and a younger age ($S_{mp}$) of the references.
Title $ARQ$ $S_{mp}$ $RQM$
@@ 5.4 Assessing the Need for Updating Reviews
Literature reviews inevitably age, yet the rate of aging varies significantly across fields (as shown in Fig. 7). The $RUI$ allows us to gauge the degree of aging of a review. Typically, fields with higher publication rates undergo faster advancements (indicated by a higher $RUI$), meaning that reviews in these areas tend to age more quickly. Conversely, in more established fields with a lower $RUI$, the aging process is correspondingly slower. Compared to relying solely on publication time to assess recency, $RUI$ provides a more accurate estimation. Examples illustrating the exploitation of the $RUI$ are provided in Tab. VII.
TABLE VII: Example Use Cases of $RUI$: $RUI$ quantifies the urgency of updating reviews across different fields.
Title Year Topic $RUI$
@@ 5.5 Integrated Use of the Proposed Indicators
In general, researchers are encouraged to apply a set of metrics, rather than relying solely on citation counts to select review articles. In practice, reviews with a $TNCSI$ close to 1 are often prioritized, as this suggests that the review is cited more frequently than most other papers in the same field, indicating widespread esteem within the field. However, supplementary indicators can provide further insights: the $IEI$ captures shifts in citation trends, the $RQM$ evaluates the quality of references, and the $RUI$ reflects how recently the review may require updating. Leveraging a diverse set of indicators allows for the efficient filtering of numerous similar reviews while mitigating the limitations inherent in single-metric citation reliance. More examples of integrated indicators usage are provided in Appendix A.5.
Although the proposed metrics provide an intuitive numerical indication, it is important to emphasize that no existing metric, including the proposed ones, can truly capture the intrinsic value of a review (particularly for systematic review). The proposed impact or quality indicators should be regarded as supplementary instruments rather than definitive measures of quality. It is our contention that the true value of a review resides in its capacity to address the specific research questions and meet the needs of its intended audience. Further explanations regarding the fair use of metrics and how to avoid their misuse are detailed in Appendix A.6.
@@ 6.1 Overview of AI-generated Literature Reviews
Traditionally, literature reviews have been manually conducted by researchers who analyze and synthesize scholarly sources to provide an overview of the current state of knowledge in a specific field. Recently, with the advancement of AI technologies, there’s been a growing interest in leveraging artificial intelligence techniques, especially LLMs, to automate or assist in the generation of the literature review [26, 114, 124, 56, 42]. Typically, users are simply required to indicate their area of research interest, and the system will then automatically generate a literature review.
@@
The automated creation of a literature review is a multidisciplinary endeavor that integrates knowledge from various fields. It relies not only on artificial intelligence technologies but also requires the merging of knowledge from other fields such as data science, bibliometrics, database engineering, etc. These components are instrumental in extracting, storing, and synthesizing relevant information from vast amounts of literature. Early attempts in AI-generated literature reviews involve training language models, e.g. LLaMA [116], on a large corpus of academic papers, research articles, and other scholarly content. These models can then be used to generate coherent and contextually relevant text based on prompts or queries related to a specific research topic. However, given that a fully trained or fine-tuned language model has no access to the latest scholarly advances, these models would not generate a literature review that includes the latest scholarly materials. Thus, most existing advanced AI-generated literature review systems contain three main steps: knowledge retrieval, synthesis, and report.
The generation of literature review begins by gathering knowledge from various sources. There is no doubt that a high-quality knowledge retrieval procedure provides a richer context for the LLMs, resulting in the production of more accurate and decent responses. The system usually employs several predefined criteria such as keywords, publication date ranges, and citation numbers to locate, filter, and retrieve relevant resources to be synthesized. These resources are not limited to academic publications, but may also include blogs, official tutorials, and GitHub repositories. Once the relevant knowledge is retrieved, the AI system aims to synthesize the gathered information to create a coherent and comprehensive literature review. This step involves analyzing the retrieved content, identifying relationships between different sources, and organizing the information into meaningful sections. The system may employ LLMs and elaborate prompts to generate the text for each section or the entire survey at once. Finally, the system seamlessly transitions into the report generation stage, where a formatted review is crafted in a cohesive and well-structured style. This stage involves converting the synthesized content into a final report which is ready for presentation. One of the simplest and most common forms of presentation is plain text. For certain systems, the generated text is arranged in various sections based on the predefined setting. Such a multi-step procedure not only streamlines the literature review generation workflow but also ensures the production of acceptable content.
Though researchers are expecting to save significant time and effort in conducting comprehensive surveys by embracing AI-generated literature reviews, numerous concerns about ethical issues and fabrications in AI-generated academic content have been raised. A critical appraisal by Zybaczynska [153] highlights that current AI systems typically fall short in providing substantial, accurate information and critical discernment. Elali et al. [29] critically examines the profound challenges brought by the fabrication and falsification of AI-generated research, underscoring its significant impact on the scientific community. Furthermore, numerous academic publishers are cautious about content created by AI. For instance, journals such as “Nature” and “Science” prohibit listing ChatGPT or other automated tools as authors on papers published in their issues. Nevertheless, we argue that the automated generation of literature reviews not intended for publication remains meaningful. It may assist researchers in staying rapidly abreast of the latest advancements in thriving fields.
@@ 6.2 Current State of AI-generated Literature Review
Most of the existing systems are only capable of generating literature reviews in plain text form and with a relatively fixed layout. A popular Github repository ChatPaper [50] retrieves papers based on a user-defined keyword with the use of the Google Scholar crawler. After a cosine similarity-based filtering of the retrieved papers, the selected papers are processed sequentially using ChatGPT. The output of Chatpaper is brief and concise, which usually contains a narrative description of the selected papers. Paper Digest [94] is an AI-based online platform that provides the service of automated review generation. Users can generate a literature review by setting specific keywords and further refining the cited source through constraints on publication dates. Similar to ChatPaper, the output of Paper Digest is also a narrative plain text description. Not all systems for generating literature reviews are fully automated. For example, Jenni AI [2] requires user interaction through “AI Command” during the review creation process, enabling the generation of highly customized literature reviews. Additionally, there are numerous online platforms [105, 14] and plugins available in the GPT Store that offer automated literature review generation services. Given the constraints on manuscript length, detailed introductions to these systems will not be provided.
Despite previous discussions indicating that certain publishers refuse to accept non-human authors, scholars have still successfully published AI-generated literature reviews in some of academic journals. Aydın et al. investigate how well can LLMs perform in the generation of the literature review. They employed well-known LLMs including ChatGPT and Google Bard to generate reviews on digital twin in healthcare [7] and meta-universes [6]. In their attempts, the Google Bard (or the ChatGPT) was first adopted to paraphrase the abstracts of papers within the last 3 years. Then, they designed several question prompts (such as “What is digital twin”) to query the LLMs and rearranged the response in a formatted style.
Although significant progress has been made in AI-generated literature reviews, there is still substantial room for improvement. Achieving automated literature reviews that match the quality of human-authored ones remains an active area of research, highlighting the need for further exploration and development. For instance, most current systems for generating literature reviews cannot extract information from tables or figures, resulting in the exclusion of these elements from the AI-generated reviews. Further details on the distinctions between AI and human-authored reviews are provided in the next subsection.
6.3 Human-authored vs. AI-generated
We first provide a descriptive analysis aimed at elucidating the distinctive approaches employed by human researchers and their AI counterparts. To prevent potential information leakage and ensure a fair analysis due to the possibility of ChatGPT having been exposed to relevant reviews during training, we selected a newly emerging field, i.e. visual prompt engineering, as the focus of our investigation. Table VIII reveals that nearly all automatic summarization systems fail to incorporate visual elements, which are highly valuable for readers. Moreover, most automated review systems lack explicit appraisal criteria, making it challenging to ensure the quality of the references included.
Review Refs Auto-Level Vis. Elems. SALSA Analysis [33]
Human [119] 160 Manullay A typically narrative, method clustering-based review aims to offer valuable insights in visual prompt learning. The selection criteria for references are not specified, but it’s clear that each reference has been thoroughly appraised.
Jenni [2] - Semi-automated $\times$ Automated searching and appraising references are not supported. The generation process relies on user interaction, and only a narrative description is provided.
AutoSurvey [124] 143 Automated $\times$ Narrative review without any visual elements. Although no formal appraisal framework is utilized, the approach incorporates semantic verification and multi-LLM evaluation. Generated reviews include narrative descriptions of key contributions, research trends, and gaps.
ChatPaper [50] 15 Automated $\times$ Retrieving relevant literature from arXiv based on multiple LLM-generated keywords without appraisal step. Each section contains plain descriptions of the related reference.
PaperDigest [94] 20 Automated $\times$ Searching papers with the user-specified keyword. It seems to appraise the quality of references with private criteria. The generated content is more like a concise summary.
askyourpdf [14] 9 Automated $\times$ No official explanations are found for how to retrieve references and appraise the quality of the literature. The generated review includes a brief analysis and description of the related concept, current state of development, and gaps.
@@
TABLE VIII: Comparisons between human-authored and AI-generated literature reviews. “Vis. Elems.” denotes the visual elements.
@@
In Fig. 8, we further present a set of scatter plots depicting the references’ quality of (a) reviews published in several academic journals and (b) reviews generated by various AI systems. To allow sufficient time for the references selected by the AI-generated review system to accumulate academic impact, a two-month interval is set between obtaining the reference list and calculating their $TNCSI$. The horizontal axis represents the reference age, and the vertical axis corresponds to the proposed $TNCSI$. The color and size of the scatter points indicate the $IEI$ of the references. A positive $IEI$ value signifies a gradually increasing citation trend, resulting in warm-colored scatter points. The larger the size of the scatter point, the greater the value of $IEI$. Conversely, when the $IEI$ value is negative, the scatter points are cool-colored, and the size decreases as the value decreases. A closer examination of Fig. 8 reveals a distinct distribution among human-authored and AI-generated depicted reviews in their respective scatter plots. Human-authored reviews exhibit clustered dots in the upper left corner, indicating a trend toward referencing more recent and influential sources (i.e., citing newly published works with limited influence at the time of writing). Conversely, AI systems predominantly cite established, high-impact papers, creating a relative sparsity in the upper left corner. We argue that this discrepancy stems from seasoned researchers’ ability to assess a paper’s value through a diverse, fine-grained, and content-based manner, rather than relying solely on citation counts. This allows them to render more precise judgments about the potential significance of a paper. In contrast, most existing AI-based review systems depend exclusively on citation counts as their primary filtering metric, limiting their capacity to identify high-quality recent publications as references. We also note the presence of a potential boundary line for the $IEI$ metric in Fig. 8, panel (b), which is likely a result of the quality control mechanisms in certain automated review systems.
@@
In summary, the utility of AI-generated reviews remains limited for several reasons. Firstly, they often lack adequate visual elements and adversely affect the overall readability and user experience. Secondly, current systems face challenges in accurately assessing the quality of newly published articles, which undermines the key advantage of AI-generated reviews—timeliness. At this stage, users are encouraged to review the selection criteria of AI-generated review systems before using them and carefully assess the newly published references included. Additionally, we would like to emphasize that researchers should fully leverage another significant advantage of AI-generated reviews: high customizability. Researchers could provide clear and specific research questions when using AI review systems to ensure the generated reviews meet their specific needs.
7 Challenges and the Future of the Literature Review

### v6

@@ 4.1 Appraising Reviews across Various Topics
A fundamental need may lie in the assessment of reviews across various fields, as most reviews are primarily aimed at readers who are relatively new to a field and may lack an in-depth understanding of the field. For instance, a researcher working in a popular field like object detection may start exploring a less familiar subfield such as optical character recognition (a subfield of computer vision). When encountering a review with only 100 citations, the researcher might dismiss its importance, unaware that in a niche area, 100 citations can actually indicate significant influence. This could lead to an oversight of critical insights in the subfield, simply due to differing citation norms between areas. The proposed $TNCSI$ offers an effective solution by automatically identifying the field of a given article and dynamically estimating its normalized impact based on the field and its current citation count. Illustrative examples are presented in Tab. 3.
Table 3: Example Use Cases of TNCSITNCSI: A similar number of citations (Cites) in various topics may lead to notably different TNCSITNCSI values.
Title Cites Topic TNCSITNCSI
@@ 4.2 The Role of Reference Quality in Selecting Reviews
As the saying goes, 'You can't make something out of nothing'. The same principle applies to literature reviews: a well-crafted review is typically built upon a thorough analysis of high-quality references. To further explore the relationship between the quality of references in literature reviews and the review's impact, we conduct a visual analysis of the distributions of $ARQ$ (representing the quality of references), $S_{MP}$ (representing the recency of references), and the corresponding $TNCSI$ (representing the scholar impact) for reviews published after 2020, as shown in Fig. 8.
Figure 8: Visualization of the Relationship between Reference Quality and TNCSITNCSI: Each point corresponds to a review paper. The TNCSITNCSI shows positive correlations with ARQARQ and $RQM$, but a negative correlation with $S_{MP}$.
Table 4: Correlation between Reference Quality and TNCSI: ``Sig.'' indicates significance, and ✓indicates the corresponding p-value less than 0.05.
Metric Pearson Pearson p-value Spearman Spearman p-value
ARQARQ 0.43 $6.99\times 10^{-40}$ 0.42 $1.11\times 10^{-37}$
$S_{mp}$ -0.27 $6.29\times 10^{-16}$ -0.32 $2.21\times 10^{-21}$
RQMRQM 0.53 $1.34\times 10^{-62}$ 0.51 $1.86\times 10^{-57}$
To further quantify the relationship between reference features and cumulative academic impact, we present the Pearson and Spearman correlation coefficients between various indicators and $TNCSI$ in Tab. 4. Both $ARQ$ and $RQM$ show moderate positive correlations with $TNCSI$, with Pearson and Spearman coefficients around 0.43 and 0.53, respectively. These correlations are highly significant, as indicated by p-values close to zero. In contrast, $S_{mp}$ exhibits a negative correlation, as indicated by Pearson and Spearman coefficients of -0.27 and -0.32, respectively. These statistically validated results demonstrate the importance of reference quality in influencing academic impact of reviews.
Considering that the quality of references exhibits a moderately strong and significant positive correlation with the scholarly impact of a review, $RQM$ can be regarded as an intuitive metric for efficiently gauging the average quality of references. In particular, as observed from Fig. 8, reviews with an $RQM$ greater than 0.8, an $ARQ$ within the range of 0.7 to 0.9, and an $S_{MP}$ not exceeding 10 are more likely to achieve higher academic impact. Illustrative examples of $RQM$ are provided in Tab. 5.
Table 5: Example Use Cases of RQMRQM: A higher RQMRQM reflects higher scholarly impact (ARQARQ) and a younger age ($S_{mp}$) of the references.
Title ARQARQ $S_{mp}$ RQMRQM
@@
4.3 Tracking Citation Momentum with IEI
Although $TNCSI$ enables cross-domain comparisons, it fails to track changes in citation trends. In such cases, the $IEI$ can offer additional insight. A higher $IEI$ suggests that the paper's rate of gaining new citations has been increasing over time, reflecting a growing level of recognition within the academic community. Conversely, an $IEI$ below zero implies that the paper’s citation rate has been declining recently. Table 6 provides examples of $IEI$ applications. In some cases, selecting a paper with a higher $IEI$ may be preferable to choosing one with a higher $TNCSI$, as it reflects a more recent consensus among researchers.
Table 6: Example Use Cases of IEIIEI: A positive IEIIEI indicates that the paper is gaining increasing attention.
Title Cites TNCSITNCSI IEIIEI
@@ 4.4 Assessing the Need for Updating Reviews
Literature reviews inevitably age, yet the rate of aging varies significantly across fields. The $RUI$ allows us to gauge the degree of aging of a review. Typically, fields with higher publication rates undergo faster advancements (indicated by a higher $RUI$), meaning that reviews in these areas tend to age more quickly. Conversely, in more established fields with a lower $RUI$, the aging process is correspondingly slower. Compared to relying solely on publication time to assess recency, $RUI$ provides a more accurate estimation. Examples illustrating the exploitation of the $RUI$ are provided in Tab. 7.
Table 7: Example Use Cases of RUIRUI: RUIRUI quantifies the urgency of updating reviews across different fields.
Title Year Topic RUIRUI
@@
4.5 Insights into Field Trends via IEI and RUI
Figure 9: Temporal Distribution of $IEI$ and $RUI$ in Surveys for Recommender Systems and Language Modeling: Higher IEIIEI and $RUI$ values in language modeling over the past two years indicate a trend of intensified development. Although surveys in popular fields tend to attract more academic attention, they also require authors to make greater efforts to keep the content up-to-date.
In the left panel of Fig. 9, we observe a rapid growth trend in the development of surveys within the field of ``language modeling'' since 2023, which is reflected by the increasing number of surveys with positive $IEI$ values. The right panel of Fig. 9 further supports this finding. Since 2023, the median $RUI$ for language modeling surveys has noticeably exceeded that of surveys in the recommendation systems field. As $RUI$ is positively correlated with the volume of publications within a field, a higher $RUI$ suggests a larger publication volume. While this trend underscores the swift progress in the language modeling domain, it also highlights the growing burden on authors to produce and maintain high-quality surveys in such a rapidly evolving and highly competitive area.
By analyzing the statistical characteristics of numerous review articles' IEI and RUI, we can uncover developmental patterns across various fields. These patterns provide valuable insights for gauging the momentum of progress within each field and advancing exploration in emerging areas.
@@ 4.6 Integrated Use of the Proposed Indicators
In general, researchers are encouraged to apply a set of metrics, rather than relying solely on citation counts to select review articles. In practice, reviews with a $TNCSI$ close to 1 are often prioritized, as this suggests that the review is cited more frequently than most other papers in the same field, indicating widespread esteem within the field. However, supplementary indicators can provide further insights: the $IEI$ captures shifts in citation trends, the $RQM$ evaluates the quality of references, and the $RUI$ reflects how recently the review may require updating. Leveraging a diverse set of indicators allows for the efficient filtering of numerous similar reviews while mitigating the limitations inherent in single-metric citation reliance.
Although the proposed metrics provide an intuitive numerical indication, it is important to emphasize that no existing metric, including the proposed ones, can truly capture the intrinsic value of a review (particularly for systematic review). The proposed impact or quality indicators should be regarded as supplementary instruments rather than definitive measures of quality. It is our contention that the true value of a review resides in its capacity to address the specific research questions and meet the needs of its intended audience. More examples of integrated indicators usage and further explanations regarding the fair use of metrics are detailed in the A.
@@ 5.1 Overview of AI-generated Literature Reviews
Traditionally, literature reviews have been manually conducted by researchers who analyze and synthesize scholarly sources to provide an overview of the current state of knowledge in a specific field. Recently, with the advancement of AI technologies, there's been a growing interest in leveraging artificial intelligence techniques, especially LLMs, to automate or assist in the generation of the literature review [118, 116, 128, 57, 43]. Typically, users are simply required to indicate their area of research interest, and the system will then automatically generate a literature review.
@@
The automated creation of a literature review is a multidisciplinary endeavor that integrates knowledge from various fields. It relies not only on artificial intelligence technologies but also requires the merging of knowledge from other fields such as data science, bibliometrics, database engineering, etc. These components are instrumental in extracting, storing, and synthesizing relevant information from vast amounts of literature. Early attempts in AI-generated literature reviews involve training language models, e.g., LLaMA [119], on a large corpus of academic papers, research articles, and other scholarly content. These models can then be used to generate coherent and contextually relevant text based on prompts or queries related to a specific research topic. However, given that a fully trained or fine-tuned language model has no access to the latest scholarly advances, these models would not generate a literature review that includes the latest scholarly materials. To overcome this limitation, most advanced AI-generated literature review systems adopt a structured pipeline consisting of four main steps: intention analysis, knowledge retrieval, synthesis, and report generation. The generation of a literature review begins with intention analysis, where the system identifies the research goals and scope. On this basis, the core process consists of knowledge retrieval and synthesis, which are closely interwoven rather than strictly sequential. Retrieval broadens the evidence base by applying criteria such as keywords, publication dates, and citation counts, and by collecting materials not only from academic publications but also from sources such as blogs, tutorials, and open-source repositories. Synthesis complements this step by analyzing the collected resources, identifying relationships among them, and organizing the information into coherent themes. By reinforcing each other, retrieval and synthesis ensure that the system produces a comprehensive and well-grounded review. Finally, the process culminates in report generation, where the integrated content is transformed into a structured and readable document.
Beyond their technical implementation, AI-generated literature reviews are often viewed as a promising tool for reducing researchers’ time and effort. However, concerns persist regarding both the reliability and the ethical implications of such content. Zybaczynska [162] observes that current AI systems frequently fail to deliver substantial, accurate information or critical judgment. Similarly, Elali et al. [30] highlight the risks of fabrication and falsification in AI-generated research, emphasizing their potentially serious consequences for the scientific community. Despite these challenges, there have been notable demonstrations of AI-assisted reviews reaching publication. For instance, Aydın et al. used large language models such as ChatGPT and Google Bard to generate reviews in areas including digital twins in healthcare [7] and the metaverse [6]. Even so, many academic publishers continue to adopt a cautious stance. Leading journals such as Nature and Science, for example, explicitly prohibit the listing of ChatGPT or other automated tools as paper authors. We contend, nevertheless, that AI-generated reviews—when not intended for direct publication—can still provide substantial value. In particular, they may serve as an efficient aid for researchers seeking to keep pace with rapid developments in fast-evolving fields.
@@ 5.2 Current State of AI-generated Literature Review
5.2.1 Comparison of Recent AI-generated Literature Review Systems
Positioned as a significant direction in AI for Research, automated review generation systems have garnered substantial attention across the research and technology communities. Early explorations—such as Paper Digest [96], Jenni AI [2], and various GPT-store plugins [106, 14]—highlight the feasibility of leveraging large language models to generate literature reviews with minimal user input. Building upon this foundation, recent efforts (see in Fig 10) have introduced more systematic and scalable approaches that emphasize structured pipelines, citation fidelity, and quality-controlled synthesis. For instance, hierarchical catalogue generation for literature review foregrounds outline construction as a first-class task, releasing a large-scale dataset and semantics-plus-structure metrics that supply supervision for catalogue-guided synthesis [161]. Building on this foundation, AutoSurvey [128] pioneers an end-to-end pipeline—seed retrieval and outline planning, subsection drafting by specialized LLMs, global integration, and iterative evaluation—demonstrating feasibility at scale. To scale capacity and preserve citation fidelity, SurveyGO [122] employs the LLM×MapReduce-V2 “convolutional stacking” architecture and uses an internal benchmark to ensure structural rigor and precise citations. Focusing on workflow quality, SurveyX [62] adopts a two-phase Preparation→Generation design that blends offline/online retrieval with an AttributeTree pre-processing pipeline and a re-polishing stage, yielding measurable gains in content quality and citation accuracy. SurveyForge [141] narrows the human–AI gap via heuristic-guided outline construction and a memory-driven scholar navigation agent (SANA), coupled with temporal- and citation-aware re-ranking to improve reference reliability and downstream drafting. Emphasizing user control and multimodality, InteractiveSurvey [131] supports outline- and section-wise RAG synthesis and produces reports enriched with structural diagrams and extracted figures/tables. From a reliability perspective, SciSage [111] introduces a citation-aware multi-agent framework that “reflects while writing,” critiquing drafts at outline, section, and document levels and re-ranking multi-source retrieval by recency and citation signals to strengthen coherence and references.
Figure 10: Comparative Overview of Recent Survey Generation Systems: The figure contrasts representative methods (HiCATGLR [161], AutoSurvey [128], SurveyX [62], SurveyForge [141], InteractiveSurvey [131], SurveyGo [122], SciSage [111]) across four key stages: intention analysis, knowledge retrieval, synthesis, and report generation. The timeline highlights their evolution from 2023 to 2025.
It is not difficult to observe that, as technologies evolve over time, automated survey systems have gradually matured across the four core components of intention analysis, knowledge retrieval, synthesis, and report generation. Compared to early explorations, many recent systems now produce reviews that are markedly more coherent, informative, and credible. The generated outputs exhibit more reasonable structural organization, richer visual elements such as extracted or synthesized figures and diagrams, and more accurate and verifiable citations. These improvements collectively enhance the readability and reliability of AI-generated reviews, allowing such systems to move beyond proof-of-concept prototypes and begin offering practical value in assisting researchers with staying abreast of emerging fields.
5.2.2 Human-authored vs. AI-generated Reviews
To highlight the differences in underlying design principles, we provide a descriptive comparison between human-authored and AI-generated reviews. Table 8 shows that nearly all automatic summarization systems fail to incorporate visual elements. Furthermore, most AI-generated review systems do not include explicit appraisal criteria, making it difficult for users to discern how quality control is enforced and thereby reinforcing the perception of these systems as black boxes.
@@ 5.2.1 Comparison of Recent AI-generated Literature Review Systems
Table 8: Comparison Between Human-authored and AI-generated Literature Reviews: "V.E." indicates the presence of visual elements.
Review Auto-Level V.E. SALSA Analysis [33]
Human [123] Manullay A typical narrative review aims to offer valuable insights in visual prompt learning. The selection criteria for references are not specified, but it's clear that each reference has been thoroughly appraised. Conducts an in-depth synthesis and offers insightful analysis.
Jenni [2] Semi-automated $\times$ Automated searching and appraising references are not supported. The generation process relies on user interaction, and only a narrative description is provided. Provides little to no synthesis or analysis.
ChatPaper [51] Automated $\times$ Retrieving relevant literature from arXiv based on multiple LLM-generated keywords without appraisal step. Each section contains plain descriptions of the related reference. Provides little to no synthesis or analysis.
PaperDigest [96] Automated $\times$ Searching papers with the user-specified keyword. It seems to appraise the quality of references with private criteria. The generated content is more like a concise summary. Provides little to no synthesis or analysis.
askyourpdf [14] Automated $\times$ No official explanations are found for how to retrieve references and appraise the quality of the literature. The generated review includes a brief analysis and description of the related concept, current state of development, and gaps. Provides little to no synthesis or analysis.
AutoSurvey [128] Automated $\times$ Narrative review without any visual elements. Although no formal appraisal framework is utilized, the approach incorporates semantic verification and multi-LLM evaluation. Generated reviews include narrative descriptions of key contributions, research trends, and gaps. Provides little to no synthesis or analysis.
SurveyGo [122] Automated No official explanations are found for how to retrieve references and appraise the quality of the literature. The generated review exhibits a clear outline structure. Offers limited synthesis and analysis.
@@
In Fig. 11, we further present a set of scatter plots depicting the references' quality of (a) reviews published in several academic journals and (b) reviews generated by various AI systems. To allow sufficient time for the references selected by the AI-generated review system to accumulate academic impact, a two-month interval is set between obtaining the reference list and calculating their $TNCSI$. The horizontal axis represents the reference age, and the vertical axis corresponds to the proposed $TNCSI$. The color and size of the scatter points indicate the $IEI$ of the references. A positive $IEI$ value signifies a gradually increasing citation trend, resulting in warm-colored scatter points. The larger the size of the scatter point, the greater the value of $IEI$. Conversely, when the $IEI$ value is negative, the scatter points are cool-colored, and the size decreases as the value decreases. A closer examination of Fig. 11 reveals a distinct distribution among human-authored and AI-generated depicted reviews in their respective scatter plots. Human-authored reviews exhibit clustered dots in the upper left corner, indicating a trend toward referencing more recent and influential sources (i.e., citing newly published works with limited influence at the time of writing). Conversely, AI systems predominantly cite established, high-impact papers, creating a relative sparsity in the upper left corner. We argue that this discrepancy stems from seasoned researchers' ability to assess a paper's value through a diverse, fine-grained, and content-based manner, rather than relying solely on citation counts. This allows them to render more precise judgments about the potential significance of a paper. In contrast, most existing AI-based review systems depend exclusively on citation counts as their primary filtering metric, limiting their capacity to identify high-quality recent publications as references. We also note the presence of a potential boundary line for the $IEI$ metric in Fig. 11, panel (b), which is likely a result of the quality control mechanisms in certain automated review systems.
@@
Overall, despite continuous technological advances, AI-generated review systems still face several critical limitations. Most systems continue to rely primarily on similarity-based retrieval, with limited attention to the quality of the retrieved literature. Although citation counts are sometimes used as a proxy, they are domain-dependent, temporally biased, and systematically disadvantage newly published work, making them an imperfect signal at best. In addition, current AI systems remain underdeveloped in terms of personalization: they seldom elicit or model users' specific research questions, and thus struggle to generate tailored surveys or guide users toward customized choices. Instead, they often imitate human-written surveys that prioritize coverage, sacrificing the customization that AI could uniquely provide. Even when visual elements are included, they are typically extracted directly from source papers rather than designed as polished, interpretive illustrations (e.g., classification diagrams, timelines, or methodological overviews). Moreover, automated approaches have yet to demonstrate the capability to robustly produce field-level artifacts such as reproducible benchmark tables or domain-specific taxonomies that systematically capture and compare the state of the art. Taken together, these limitations suggest a growing need to explore more reliable, transparent, and intent-aligned approaches to AI-generated literature reviews.
5.3 Toward Reliable, Transparent, and Intent-Aligned AI-generated Reviews
Although the capabilities of AI-generated review systems are steadily improving, their adoption in practice remains modest. This may be partly due to ongoing challenges related to reliability, transparency, and the lack of intent-aware customization.
A key factor limiting the reliability of current systems lies in the lack of mechanisms for evaluating the credibility and quality of referenced literature. Many models retrieve papers based solely on lexical or embedding similarity, without regard to publication venue, peer-review status, or citation history—signals that are often crucial in scholarly contexts. Enhancing reliability may require the integration of explicit citation-quality management components that filter or re-rank literature based on such criteria [133, 156]. Furthermore, deeper exploitation of LLMs’ reasoning capabilities [37] could allow for more substantive synthesis across sources, moving beyond surface-level summaries toward analytical, comparative, and even critical review writing. Emerging techniques, such as diffusion-based scientific visualization [22], may also contribute to reliability by enabling the automated generation of polished, interpretable visual elements.
Transparency remains a persistent challenge. Most existing systems function as black boxes: users are seldom informed about how references are retrieved, filtered, or retained. The lack of interpretable intermediate steps or explainable selection criteria undermines trust and limits rigorous evaluation. In contrast, greater transparency in AI-generated reviews can markedly strengthen users’ confidence, as it enables them to better understand and assess the reliability of the outputs. To foster both trust and accountability, AI review systems should disclose their retrieval logic, inclusion and exclusion criteria, and synthesis provenance in a manner that is accessible to both technical and non-technical audiences. Such transparent pipelines not only facilitate critical appraisal but also make it easier to identify errors, biases, and omissions in the generated content.
Beyond reliability and transparency, one of the most overlooked opportunities of AI-generated reviews may lie in intent-aligned customization. Unlike human-authored reviews—which, once published, must serve a broad and heterogeneous readership—AI systems can adapt outputs dynamically to specific user goals. Depending on the research purpose, a system could generate a concise state-of-the-art overview for practitioners, a mapping review to highlight gaps in the literature, or a full systematic review to support meta-analysis. By tailoring outputs to user intent, such systems can deliver reviews that are more concise, targeted, and practically useful, with flexibility in analytic depth, coverage, and timeliness. This shift toward purpose-driven review generation underscores a unique advantage of AI: not simply replicating established formats, but reimagining the practice of literature reviewing around individual needs and research contexts.
6 Challenges and the Future of the Literature Review in PAMI